OpenStudy (anonymous):

If f(x+1)-f(x)=2 and f(1)=5 what is f(6)

OpenStudy (anonymous):

Find x using f(1)=5 using x find f(6)

OpenStudy (anonymous):

\[f(x+1)-f(x) = 2\] \[f(1+1)-f(1)=2\]\[f(2)-5=2 \ \ \rightarrow f(2)=7\] Continue plugging in and you'll find this pattern: 5, 7, 9, 11, 13, 15, ...